People in the aerospace industry believe the cost of a space project is a function of the mass of the major object being sent into space. Use the following data to develop a regression model to predict the cost of a space project by the mass of the space object. Determine r2 and se. Weight (tons) Cost ($ millions) 1.897 $ 53.6 3.019 184.6 0.453 6.4 0.996 23.5 1.058 32.9 2.100 110.4 2.385 104.6

ŷ = enter a number rounded to 4 decimal places  * + enter a number rounded to 4 decimal places * x
r2 = enter a number rounded to 3 decimal places  **
se = enter a number rounded to 3 decimal places  **
Answer #1

1)

y^ =-39.3167+66.4442x

2)

SST=Syy= 23,718.4886
SSE =Syy-(Sxy)2/Sxx= 2,215.742
SSR =(Sxy)2/Sxx = 21,502.7465
Coeffficient of determination R^2 =SSR/SST= 0.907

3)

s2 =SSE/(n-2)= 443.1484
std error σ              = =se =√s2= 21.051